Question:A bar of copper of length 0.75 m and a bar of a staniless steel of length 1.25 m are joined end to end,each having a cross-sectional area of 10^-4 m^2 .The free ends of the copper and the steel bars are maintained at 100 degree C and 0 degree C respectively.1)Find the temperature of the copper-steel junction   .2)Calculate the rate of heat transmission  across the junction.Thermal conductivities of copper and steel  are400  and 14 joule+sec.m.kelvin respectively .

Grade:12th pass

1 Answers

Eshan
5 years ago
Since the rods are connected in series, the rate of heat supplied would be constant. Let the temperature at the junction of the rod beT.

Therefore,\dfrac{dH}{dt}=\dfrac{K_1A(100-T)}{0.75}=\dfrac{K_2AT}{1.25}

Solving the above equation will give the temperature as well as the rate of heat transmission.